OpenStudy (anonymous):

Please help me multiply this rational expression: -4s/9t * -9s^2/2

OpenStudy (anonymous):

9 and 9 cancel out

OpenStudy (anonymous):

The arcsin(sqrt(2)/2) = pi/4 and 3pi/4

OpenStudy (anonymous):

so you have left \[\frac{ (4s)(s ^{2}) }{ 2 }\]

OpenStudy (anonymous):

which is the same as \[\frac{ 4s ^{3} }{ 2t }\]

OpenStudy (anonymous):

i forgot the t in the first one

OpenStudy (anonymous):

but if you simplify you should get \[\frac{ 2s ^{3} }{ }\]

OpenStudy (anonymous):

\[\frac{ 2s ^{3} }{ t }\]

OpenStudy (anonymous):

Thank you @yomylola!